# Feature Alignment: Rethinking Efficient Active Learning via Proxy in the Context of Pre-trained Models

### v1

@@ 6.5 Ablation Study
Our ablation study was conducted on CIFAR-10, CIFAR-100, and Pets and the results are presented in table 3. Additional ablation studies on other active learning strategies are provided in appendix 11. Across all three datasets, updating pre-computed features (feature alignment) consistently led to a noticeable performance boost. Meanwhile, switching training methods (training alignment) improved performance on all datasets except for the Pets dataset, suggesting that our method may not precisely determine the optimal point for switching training methods. However, considering the practical scenario where we cannot pre-determine whether using FT or LP-FT for training is better, our approach still aids in selecting a preferable training method.
Feature Training CIFAR-10 CIFAR-100 Pets
alignment alignment
@@
Table 3: Ablation Study. Comparison of the average sample saving ratio of solely updating pre-computed features (training the final model using fine-tuning), solely switching training methods (using pre-trained features), and the complete ASVP approach (both updating pre-computed features and switching training methods) under different datasets with margin sampling.
@@ 6.6 Influence from Position of Updating Features
We examined the impact of the position of updating pre-computed features on CIFAR-10. The number of saved samples throughout the entire active learning iteration process is depicted in fig. 8, with the average savings ratio presented in table 4. The choice of updating pre-computed features does indeed affect the final performance of ASVP. Nevertheless, across the six different positions explored in the experiments, ASVP consistently outperforms the method relying solely on pre-training features (SVPp). Furthermore, our proposed method for determining the update location based on LogME-PED proves to be effective. On CIFAR-10, this method selects the update of pre-computed features at 600 labels, ranking as the second-best among the six experimental positions, as shown in table 4. The positions of updating pre-computed features estimated through LogME-PED are outlined in appendix 12.
Change Saving Change Saving
at Ratio at Ratio
@@
Figure 8: The impact of the position of updating pre-computed features on the savings of label amounts in CIFAR-10 across each active learning iteration. The equivalent non-active learning label amount refers to the number of samples selected by a random baseline to achieve the same accuracy as active learning
6.7 Improvement from Small Batchsize
The training costs associated with deep neural networks often lead practitioners to augment the number of samples selected per active learning iteration to ensure computational feasibility. Unfortunately, this strategy tends to result in selecting some redundant samples. Given the efficiency of our method, we can alleviate this problem. We conducted experiments on ImageNet, randomly selecting an initial pool of 10,000 samples. Subsequently, we performed 9, 45, 180, and 900 active learning iterations using margin sampling, selecting a total of 100,000 samples. The accuracy of the final model and the total sample selection time for both FT and LP-FT training are presented in table 5. The results show a general improvement in active learning performance with an increase in the number of iterations. However, beyond 180 iterations, the performance improvement diminishes, indicating a limit to the benefits of increasing the number of active learning iterations.
# AL Training Method Sampling
iterations FT LP-FT Time / hrs
9 16.64$\pm$2.06 30.22$\pm$1.85 1.1
45 17.66$\pm$1.37 31.20$\pm$0.40 1.5
180 19.43$\pm$0.86 32.00$\pm$0.46 3.7
900 18.21$\pm$0.67 31.64$\pm$0.24 15.8
@@ 7 Conclusion
Active learning on pre-trained models shows label efficiency but demands significant computational time, whereas current efficient active learning methods tend to compromise a notable fraction of active learning performance, resulting in increased overall costs. To address this challenge, this paper attributes the reduced performance of SVPp to (1) the absence of samples necessary for preserving valuable pre-trained information and (2) an increased selection of redundant samples due to fine-tuned features becoming better than pre-trained features as the annotation increases. Building upon this analysis, we propose ASVP, a two-fold alignment approach to improve the effectiveness of the efficient active learning method with similar or marginally increased computational time. Additionally, we introduce the sample savings ratio as a metric to assess the effectiveness of efficient active learning, providing a straightforward measure for labeling cost savings. Experiments show that our proposed ASVP saves similar or greater total costs in most cases, while still maintaining computational efficiency.
Supplementary Material
8 Hyperparameters
@@
The hyper-parameters of fine-tuning are shown in table 6. For ImageNet, we adopted the fine-tuning parameters following the prior work [4]. Setting the learning rates separately for the classifier head and the backbone. For CIFAR-10, CIFAR-100, and Pets, we followed prior work [4] to set the learning rate of the classifier as 0.1 and to search other hyper-parameters. The hyper-parameters are detailed in table 6. The learning rate of the backbone was searched from (0.01, 0.05, 0.1 0.3) and the weight decay from (0, 0.0001). Additionally, to maintain consistency in training iterations between fine-tuning and LP-FT, we set the number of fine-tuning epochs as the sum of linear probing and fine-tuning epochs in LP-FT training.
@@
Table 7: The hyper-parameters of the linear-probing then fine-tuning.
@@
9 Accuracy and Sample Savings
(a) Accuracy
(b) Sample Savings
Figure 9: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the ImageNet, where active learning strategy is BADGE.
(a) Accuracy
(b) Sample Savings
Figure 10: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the ImageNet, where active learning strategy is confidence.
(a) Accuracy
(b) Sample Savings
Figure 11: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the CIFAR-10, where active learning strategy is margin.
(a) Accuracy
(b) Sample Savings
Figure 12: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the CIFAR-10, where active learning strategy is BADGE.
(a) Accuracy
(b) Sample Savings
Figure 13: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the CIFAR-10, where active learning strategy is confidence.
(a) Accuracy
(b) Sample Savings
Figure 14: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the CIFAR-10, where active learning strategy is activeft(al).
(a) Accuracy
(b) Sample Savings
Figure 15: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the CIFAR-100, where active learning strategy is margin.
(a) Accuracy
(b) Sample Savings
Figure 16: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the CIFAR-100, where active learning strategy is BADGE.
(a) Accuracy
(b) Sample Savings
Figure 17: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the CIFAR-100, where active learning strategy is confidence.
(a) Accuracy
(b) Sample Savings
Figure 18: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the CIFAR-100, where active learning strategy is activeft(al).
(a) Accuracy
(b) Sample Savings
Figure 19: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the Pets, where active learning strategy is margin.
(a) Accuracy
(b) Sample Savings
Figure 20: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the Pets, where active learning strategy is BADGE.
(a) Accuracy
(b) Sample Savings
Figure 21: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the Pets, where active learning strategy is confidence.
(a) Accuracy
(b) Sample Savings
Figure 22: The accuracy and equivalent non-active learning label amounts of the standard method, SVPp and our method ASVP on the Pets, where active learning strategy is activeft(al).
10 Computation Time
@@
CIFAR-10 CIFAR-100 Pets
Standard 1.4 hrs 5.5 hrs 1.5 hrs
SVPp 0.07 hrs 0.08 hrs 0.03 hrs
ASVP 0.14 hrs 0.14 hrs 0.05 hrs
@@
11 Ablation Study
@@
Feature Training CIFAR-10 CIFAR-100 Pets
alignment alignment
@@
Feature Training CIFAR-10 CIFAR-100 Pets
alignment alignment
@@
Feature Training CIFAR-10 CIFAR-100 Pets
alignment alignment
@@
12 Postion of Updating pre-computed features
The ASVP method uses the difference in LogME-PED scores between consecutive active learning iterations to decide if updating pre-computed features is required. The absolute differences in LogME-PED scores are illustrated in fig. 23, fig. 24, fig. 25. When the threshold is defined as 1, the updating positions are detailed in table 12.
@@ 12 Postion of Updating pre-computed features
ImageNet No $\geq$ 100k
CIFAR-10 Yes 600
CIFAR-100 Yes 2000
@@
Figure 23: The absolute differences in LogME-PED scores between consecutive active learning iterations on CIFAR-10.
Figure 24: The absolute differences in LogME-PED scores between consecutive active learning iterations on CIFAR-100.
Figure 25: The absolute differences in LogME-PED scores between consecutive active learning iterations on Pets.
13 Computation Efficiency and Overall Cost
@@
In fig. 29-fig. 37, we illustrate the overall costs (the labeling cost and the active learning training cost) and computational efficiency achieved by the standard active learning method, SVPp, and our proposed ASVP across different datasets. Compared to the existing efficient active learning method, SVPp, our method significantly improves the overall cost while maintaining the computational efficiency. ASVP achieves comparable overall costs while significantly reducing active learning computation time compared to the standard active learning method. This effectively addresses the trade-off practitioners encounter between computational efficiency and total costs, offering confidence in the adoption of efficient active learning methods.
@@
Figure 26: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on ImageNet using margin sampling.
Figure 27: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on ImageNet using BADGE.
Figure 28: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on ImageNet using confidence sampling.
@@
Figure 29: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on CIFAR-10 using margin sampling.
Figure 30: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on CIFAR-10 using BADGE.
@@
Figure 31: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on CIFAR-10 using confidence sampling.
Figure 32: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on CIFAR-100 using margin sampling.
Figure 33: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on CIFAR-100 using BADGE.
Figure 34: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on CIFAR-100 using confidence sampling.
@@
Figure 35: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on Pets using margin sampling.
Figure 36: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on Pets using BADGE.
@@
Figure 37: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on Pets using confidence sampling.

### v2

@@ 6.5 Ablation Study
Our ablation study was conducted on CIFAR-10, CIFAR-100, and Pets, using 10, 15, and 10 active learning iterations, respectively, with each iteration querying 200, 400, and 200 samples. All other settings follow those described in sec. 6.1. The results with margin sampling are shown in Table 3, while additional ablation studies with other active learning strategies are provided in Appendix C. Across all three datasets, updating pre-computed features consistently led to a noticeable performance boost. Meanwhile, switching training methods improved performance on all datasets except for the Pets dataset, suggesting that our method may not precisely determine the optimal point for switching training methods. However, considering the practical scenario where we cannot pre-determine whether using FT or LP-FT for training is better, our approach still aids in selecting a preferable training method.
Table 3: Ablation Study. Comparison of the average sample saving ratio of only updating pre-computed features (training the final model using fine-tuning), solely switching training methods (using pre-trained features), and the complete ASVP approach (both updating pre-computed features and switching training methods) under different datasets with margin sampling.
Update Pre-computed Features Switch Training Methods CIFAR-10 CIFAR-100 Pets
@@ 6.6 Influence from Position of Updating Features
We examined the impact of the position of updating pre-computed features on CIFAR-10. The average savings ratio is presented in table 4. The choice of updating pre-computed features does indeed affect the final performance of ASVP. Nevertheless, across the six different positions explored in the experiments, ASVP consistently outperforms the method relying solely on pre-training features (SVPp). Furthermore, our proposed method for determining the update location based on LogME-PED proves to be effective. On CIFAR-10, this method selects the update of pre-computed features at 600 labels, ranking as the second-best among the six experimental positions, as shown in table 4. The positions of updating pre-computed features estimated through LogME-PED are outlined in appendix D.
Change at Avg. Sample Saving Ratio Change at Avg. Sample Saving Ratio
@@ 7 Conclusion
Active learning on pre-trained models offers label efficiency but requires significant computational time, whereas current efficient active learning methods tend to compromise a notable fraction of active learning performance, resulting in increased overall costs. To address this challenge, this paper attributes the reduced performance of SVPp to (1) the absence of samples necessary to preserve fine-tuned features similar to the pre-trained features and (2) an increased selection of redundant samples due to fine-tuned features becoming better than pre-trained features as the annotation increases. Building upon this analysis, we propose a simple and effective method, ASVP, to improve the effectiveness (in terms of accuracy) of efficient active learning with comparable or marginally increased computational time. Additionally, we introduce the sample savings ratio as a metric to assess the effectiveness of efficient active learning, providing a straightforward measure for labeling cost savings. Experiments show that our proposed ASVP saves similar or greater total costs in most cases while maintaining computational efficiency.
8 Limitations and Future Directions
Due to limited computational resources, this paper evaluates efficient active learning methods’ labeling and training costs using a ResNet-50 model with a specific pre-training method, BYOL-EMAN (Grill et al., 2020; Cai et al., 2021). Although ResNet-50 serves as a practical choice for our experiments and provides a reliable baseline, larger models achieve higher accuracy and are often more helpful in practical applications (Zhai et al., 2022). Existing research shows that larger models exhibit better label efficiency, requiring fewer labeled samples to reach the same accuracy as smaller models (Chen et al., 2020b). This implies that in active learning, using larger models may reduce labeling costs while increasing training costs. Therefore, the total cost savings achieved by both standard and efficient active learning methods are closely tied to model size. Evaluating the impact of model size on total cost savings, and exploring efficient active learning methods that can effectively reduce sample selection time and total cost (labeling and training costs) in the context of larger models, are important directions for future research.
Furthermore, when transferring to downstream datasets (used for active learning), different pre-trained models exhibit varying performance gaps between fine-tuning and linear probing. A smaller gap suggests higher-quality pre-trained features, making it easier for proxy models based on these features to select samples similar to those chosen by fine-tuned models. This means that proxy model-based active learning may compromise less on performance compared to standard active learning. Therefore, exploring how to select suitable pre-trained models is another interesting direction for future research.
Appendix A Hyperparameters
@@
The hyper-parameters of fine-tuning are shown in table 5. For ImageNet, we adopted the fine-tuning parameters following the prior work (Cai et al., 2021). Setting the learning rates separately for the classifier head and the backbone. For CIFAR-10, CIFAR-100, and Pets, we followed prior work (Cai et al., 2021) to set the learning rate of the classifier as 0.1 and to search other hyper-parameters. The hyper-parameters are detailed in table 5. The learning rate of the backbone was searched from (0.01, 0.05, 0.1 0.3) and the weight decay from (0, 0.0001). Additionally, to maintain consistency in training iterations between fine-tuning and LP-FT, we set the number of fine-tuning epochs as the sum of linear probing and fine-tuning epochs in LP-FT training.
Table 5: The hyper-parameters of the fine-tuning.
@@
Appendix B Computation Time
@@
CIFAR-10 (hrs) CIFAR-100 (hrs) Pets (hrs)
Standard 6.94 20.96 1.50
SVPp 0.22 0.13 0.03
ASVP 0.85 1.10 0.05
Appendix C Ablation Study
@@
Feature alignment Training alignment CIFAR-10 CIFAR-100 Pets
@@
Feature alignment Training alignment CIFAR-10 CIFAR-100 Pets
@@
Feature alignment Training alignment CIFAR-10 CIFAR-100 Pets
@@
Appendix D Postion of Updating Pre-computed Features
The ASVP method uses the difference in LogME-PED scores between consecutive active learning iterations to decide if updating pre-computed features is required. The absolute differences in LogME-PED scores are illustrated in fig. 14(a), fig. 14(b), fig. 14(c). When the threshold is defined as 1, the updating positions are detailed in table 11.
@@ Appendix D Postion of Updating Pre-computed Features
ImageNet Yes 150k
CIFAR-10 Yes 600 and 2000
CIFAR-100 Yes 2000 and 6000
@@
Figure 14: The absolute differences in LogME-PED scores between consecutive active learning iterations on (a) CIFAR-10, (b) CIFAR-100 and (c) Pets.
Appendix E Results of other Active Learning Strategies
@@
In fig. 23-fig. 26, we illustrate the overall costs (the labeling cost and the active learning training cost) and computational efficiency achieved by the standard active learning method, SVPp, and our proposed ASVP across different datasets. Compared to the existing efficient active learning method, SVPp, our method significantly improves the overall cost while maintaining the computational efficiency. ASVP achieves comparable overall costs while significantly reducing active learning computation time compared to the standard active learning method. This effectively addresses the trade-off practitioners encounter between computational efficiency and total costs, offering confidence in adopting efficient active learning methods.
(a) Margin
(b) BADGE
(c) Confidence
@@
Figure 15: The accuracy of the standard method, SVPp, and our method ASVP on the ImageNet using (a) Margin, (b) BADGE and (c) Confidence sampling.
(a) Margin
(b) BADGE
(c) Confidence
Figure 16: The equivalent number of non-active learning label amounts comparison of the standard method, SVPp, and our method ASVP on ImageNet using (a) Margin, (b) BADGE and (c) Confidence sampling. The equivalent non-active learning label amount refers to the number of samples selected by a random baseline to achieve the same accuracy as active learning.
(a) Margin
(b) BADGE
(c) Confidence
(d) ActiveFT(al)
(e) Coreset
@@
Figure 17: The accuracy of the standard method, SVPp, and our method ASVP on the CIFAR-10 using (a) Margin, (b) BADGE, (c) Confidence, (d) ActiveFT(al) and (e) Coreset sampling.
(a) Margin
(b) BADGE
(c) Confidence
(d) ActiveFT(al)
(e) Coreset
Figure 18: The equivalent number of non-active learning label amounts comparison of the standard method, SVPp, and our method ASVP on the CIFAR-10 using (a) Margin, (b) BADGE, (c) Confidence, (d) ActiveFT(al) and (e) Coreset sampling. The equivalent non-active learning label amount refers to the number of samples selected by a random baseline to achieve the same accuracy as active learning.
(a) Margin
(b) BADGE
(c) Confidence
(d) ActiveFT(al)
(e) Coreset
@@
Figure 19: The accuracy of the standard method, SVPp, and our method ASVP on the CIFAR-100 using (a) Margin, (b) BADGE, (c) Confidence, (d) ActiveFT(al) and (e) Coreset sampling.
(a) Margin
(b) BADGE
(c) Confidence
(d) ActiveFT(al)
(e) Coreset
Figure 20: The equivalent number of non-active learning label amounts comparison of the standard method, SVPp, and our method ASVP on the CIFAR-100 using (a) Margin, (b) BADGE, (c) Confidence, (d) ActiveFT(al) and (e) Coreset sampling. The equivalent non-active learning label amount refers to the number of samples selected by a random baseline to achieve the same accuracy as active learning.
(a) Margin
(b) BADGE
(c) Confidence
(d) ActiveFT(al)
(e) Coreset
@@
Figure 21: The accuracy of the standard method, SVPp, and our method ASVP on the Pets using (a) Margin, (b) BADGE, (c) Confidence, (d) ActiveFT(al) and (e) Coreset sampling.
(a) Margin
(b) BADGE
(c) Confidence
(d) ActiveFT(al)
(e) Coreset
Figure 22: The equivalent number of non-active learning label amounts comparison of the standard method, SVPp, and our method ASVP on the Pets using (a) Margin, (b) BADGE, (c) Confidence, (d) ActiveFT(al) and (e) Coreset sampling. The equivalent non-active learning label amount refers to the number of samples selected by a random baseline to achieve the same accuracy as active learning.
(a) Margin
(b) BADGE
(c) Confidence
@@
Figure 23: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on ImageNet using (a) Margin, (b) BADGE and (c) Confidence sampling.
(a) Margin
(b) BADGE
(c) Confidence
(d) Coreset
(e) ActiveFT (al)
Figure 24: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on CIFAR-10 using (a) Margin, (b) BADGE, (c) Confidence, (d) Coreset and (e) ActiveFT (al) sampling.
(a) Margin
(b) BADGE
(c) Confidence
(d) Coreset
(e) ActiveFT (al)
Figure 25: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on CIFAR-100 using (a) Margin, (b) BADGE and (c) Confidence sampling.
(a) Margin
(b) BADGE
(c) Confidence
(d) Coreset
(e) ActiveFT (al)
Figure 26: Computation efficiency and overall cost comparison of the standard active learning method, SVPp, and our method ASVP on Pets using (a) Margin, (b) BADGE and (c) Confidence sampling.
Appendix F Improvement from Small Batchsize
The training costs associated with deep neural networks often lead practitioners to increase the number of samples selected per active learning iteration to ensure computational feasibility. Unfortunately, this strategy tends to result in selecting some redundant samples. Given the efficiency of our method, we can alleviate this problem. We conducted experiments on ImageNet, randomly selecting an initial pool of 10,000 samples. Subsequently, we performed 9, 45, 180, and 900 active learning iterations using margin sampling, selecting a total of 100,000 samples. The accuracy of the final model and the total sample selection time for both FT and LP-FT training are presented in table 12. The results show a general improvement in active learning performance with an increase in the number of iterations. However, beyond 180 iterations, the performance improvement diminishes, indicating a limit to the benefits of increasing the number of active learning iterations.
# AL Training Method Sampling
iterations FT LP-FT Time (hrs)
9 16.64$\pm$2.06 30.22$\pm$1.85 1.1
45 17.66$\pm$1.37 31.20$\pm$0.40 1.5
180 19.43$\pm$0.86 32.00$\pm$0.46 3.7
900 18.21$\pm$0.67 31.64$\pm$0.24 15.8
Appendix G Replacement Experiment on CIFAR-10
We conducted replacement experiments on CIFAR-10 to demonstrate the impact of replacing missing samples from regions A1, A2, B1, and B2 on SVPp active learning performance. As shown in Figure 27, similar to the ImageNet replacement experiments, replacing samples from regions A1 and B1 results in little change in active learning performance, whereas replacing samples from regions A2 and B2 improves performance.
We also used Dataset Maps to illustrate the difficulty distribution of samples from different regions. The Dataset Map calculates the average and standard deviation of the model’s confidence for the correct class (need true labels) using training dynamics from the entire dataset. Due to the simplicity of CIFAR-10, the Dataset Map shows very high mean confidence for almost all samples. Therefore, the difficulty based on mean confidence, as used in Section 4.1, would categorize nearly all samples as “easy”.
Instead, we classified sample difficulty based on the standard deviation of the model’s confidence in the correct class. Specifically, we defined the difficulty levels as follows: very hard (std. >= 0.15), hard (std. >= 0.1 and std. < 0.15), medium (std. >= 0.05 and std. < 0.10), and easy (std. < 0.05). The resulting distribution of different difficulty levels across regions is shown in Figure 28, where the proportion of each difficulty level within the entire dataset is displayed in the last column.
Figure 27: The impact of different regions on SVPp active learning performance. Replacing samples selected by the proxy model with those from regions A1, A2, B1, B2. The active learning gain refers to the difference in accuracy between active learning and the random baseline on CIFAR-10.
(a) A1
(b) A2
(c) B1
(d) B2
Figure 28: Difficulty distribution of samples from different regions on CIFAR-10.